### v1

# WHAT-IF: Exploring Branching Narratives by Meta-Prompting Large Language Models

Abstract
WHAT-IF—Writing a Hero’s Alternate Timeline through Interactive Fiction—is a system that uses zero-shot meta-prompting to create branching narratives from a prewritten story. Played as an interactive fiction (IF) game, WHAT-IF lets the player choose between decisions that the large language model (LLM) GPT-4 generates as possible branches in the story. Starting with an existing linear plot as input, a branch is created at each key decision taken by the main character. By meta-prompting the LLM to consider the major plot points from the story, the system produces coherent and well-structured alternate storylines. WHAT-IF stores the branching plot tree in a graph which helps it to both keep track of the story for prompting and maintain the structure for the final IF system. A video demo of our system can be found here: https://youtu.be/8vBqjqtupcc.
WHAT-IF: Exploring Branching Narratives by Meta-Prompting Large Language Models
Runsheng "Anson" Huang University of Pennsylvania rhuang99@seas.upenn.edu Lara J. Martin University of Maryland, Baltimore County laramar@umbc.edu Chris Callison-Burch University of Pennsylvania ccb@seas.upenn.edu
1 Introduction
Interactive fiction (IF) is a form of text-based story where a user performs some action that moves the plot along. From Choose Your Own Adventure™ books to open-world video games, branching narratives are an interesting element of interactive fiction that enables the reader to change the outcome of the narrative.
In this paper, we present WHAT-IF (Writing a Hero’s Alternate Timeline through Interactive Fiction)111https://what-if-game.github.io/, a system for controllable, branching narrative generation in English. Starting with a human-written plot, WHAT-IF produces distinct changes in the plot by considering alternate actions the main character could have taken. We create these choice points by having a large language model (LLM) generate a new sub-goal for the character, which significantly changes subsequent plot points in the story. We use this LLM-driven branching narrative to create an interactive fiction system (Fig. 1) allowing the user to choose which branches to follow in the story tree.
In contrast with previous LLM-driven interactive fiction systems like AI Dungeon222https://aidungeon.com/, which rely entirely on the user to move the story along, WHAT-IF generates more constrained stories where the user has a degree of control but the stories generated by the LLM maintain a high degree of thematic consistency with the original story, even with many possible endings.
Structure can help LLMs better understand stories Dong et al. (2023). WHAT-IF breaks down the narrative into events so that the structure of the plot can be tracked throughout the generation. A branching plot structure, such as that shown in Figure 2, is generated before the user plays the game. Each branch of the story will lead to an ending.
As each story requires asking different questions to generate hypothetical alternate plots, we also employ meta-prompting. That is, the system generates story-specific prompts to be fed into another instance of the LLM to answer.
Although we present the system as an interactive fiction game, we believe that WHAT-IF could be used to help authors explore alternative paths in storylines they are writing. See Swanson et al. (2021); Yuan et al. (2022); Zhu et al. (2023a); Gero et al. (2023) for some examples of writing aids using LLMs. Park et al. (2023) even look at using ChatGPT to help author IFs.
In the rest of this paper, we will discuss related work about using LLMs for interactive fiction and story generation, describe the architecture of WHAT-IF, and then give example narratives that were generated.
2 Related Work: Storytelling with LLMs
Story generation has a long history in classical planning Meehan (1977); Lebowitz (1984); Young et al. (2013), where the resulting story comes from the plan extracted from a graph of plot points. These systems enable long-term coherent story plots to form, but have no capabilities for text generation.
On the other hand, large language models (LLMs) yield flexible, grammatical natural language generation. While the previous generation of LLMs (GPT-2 era) had mixed results with story generation See et al. (2019), modern LLMs are telling much better stories and are almost generating at a human level Xie et al. (2023).
There has been plenty of work on controlled story generation with LLMs Lal et al. (2021); Razumovskaia et al. (2024); Yang et al. (2022); Qin and Zhao (2022); Zhu et al. (2023b). Most of this work consists of higher-level “planning” with LLMs, where an outline or set of keywords is expanded into a longer-form story. There have also been recent works using retrieval-augmented LLMs for story generation Zhu et al. (2023a); Wen et al. (2023). However, people have only just begun to explore mixing more traditional AI planning techniques with LLMs, namely by generating components that would be fed into a planner or using pre-conditions and effects in a similar manner to a planner Martin (2021); Simon and Muise (2022); Kelly et al. (2023); Ye et al. (2023).
LLMs have also been used for story-generation–related tasks, including generating story worlds Johnson-Bey et al. (2023) and doing story sifting Méndez and Gervás (2023)—that is, creating a story from a collection of facts.
2.1 Interactive Narrative Generation
In line with story generation’s planning history, there have been plenty of interactive narrative systems using planning methods Riedl and Bulitko (2013); Sanghrajka (2019); Ware and Siler (2021). So far, neural IF work has included tasks such as generating interactive fiction worlds Ammanabrolu et al. (2020), scenes Kumaran et al. (2023), branching quests De Lima et al. (2021), or choices for the user Harmon and Rutman (2023). There has been a collection of work looking at interactive fiction through roleplaying as well, using recurrent language models Louis and Sutton (2018), reinforcement learning Martin et al. (2018); Ammanabrolu et al. (2021), and LLMs Callison-Burch et al. (2022); Ashby et al. (2023); Wang et al. (2024); Shao et al. (2023); Yan et al. (2023).
There are a couple of systems we’d like to point out in particular to compare to our system WHAT-IF. Pei et al. (2024)’s SWAG story generation system uses an LLM to both generate actions (story events) and then select the best action using Direct Preference Optimization (DPO) Rafailov et al. (2023). Although we are not selecting the action—the user is, we are relying on the LLM to generate story branches. However, instead of finetuning on pre-written stories, WHAT-IF uses zero-shot prompting to generate new events.
Similarly, like Harmon and Rutman (2023)’s works, we use LLMs to generate branching narratives by exploring character choices. Our system expands on this work by, instead of using story-specific questions as part of the prompt, WHAT-IF uses meta-prompts so that the system can ask itself story-specific questions to aid in generation.
3 System Architecture
WHAT-IF (Fig. 3) consists of five major phases: initializing the branching plot tree (§3.1), extracting key events (§3.2), generating meta-prompts (§3.3), branching & merging the tree (§3.4), and narrating the game (§3.5). These components ensure the system chooses logical branching points, maintains a coherent and consistent story structure in alternate storylines, and creates an immersive playing experience.
3.1 Initializing the Branching Plot Tree
Most interactive fiction games tell the story of the protagonist (the player) embarking on a journey full of challenges, making key decisions, and reaching their goals along the way. The major challenge for human interactive fiction writers when designing branching narratives is to propose meaningful key decisions that offer players agency in shaping the story. Inspired by De Lima et al. (2021)’s previous work of adopting a quest tree in story planning and branching, we create a branching plot tree structure to capture key decisions in a given story plot as branching points. Here, we define a key decision as the decision a main character takes in response to a given state (obstacles, dangers, etc.) which influences how the rest of the story unfolds.
The tree is constructed as follows (illustrated in Figure 4):
Each node contains the current state () and goal () of the character, the key decision () taken in the original storyline, and the alternate decision () given the same state and goal that would lead to an alternate storyline.
Each edge contains a list of three sentences corresponding to events events that lead the character to move from one node to another: (1) the decision made by the character— or as full sentences , (2) the event resulting from the decision & leading to the next state, and (3) the next state of the character . We will refer to these as edge events.
The state represents the current circumstances of the character after experiencing the previous events . The goal is what the character tries to accomplish at the given state (defeat monster, save villagers, etc.) The key decision is the decision taken in the original storyline and the alternate decision is a hypothetical decision that the character could have taken to achieve . Either decision directly leads to events of the corresponding edge to the next node.
In this phase (#1 & #5 in Figure 3), we prompt GPT-4 to summarize the story into a storyline in the branching plot tree and imagine the alternate decision at each node. A storyline is defined as a linear traversal of the tree from the root node to one of the leaf nodes. The strong causal relationship among these variables ensures it can capture meaningful decisions, sometimes hidden, in a plot that propels the story forward. Note that extracting events from all edges of a storyline would give us the full plot summary.
3.2 Extracting Key Events
In our exploratory experiments, we observed that when prompted to generate alternate storylines, GPT-4 tends to imagine stories that either deviate too much from the original story or delve into specific details which leads to story stagnation. One effective solution is providing a story structure.
A well-known story structure is the Three-act Structure that divides the story into three phases: Setup, Confrontation, and Resolution Field (1979). The backbone of this structure is three major plot points: Inciting Incident, Crisis, and Climax. These major plot points force the protagonist to experience and grow and drive the tension up and down, leading to an engaging story.
Therefore, we prompt GPT-4 to identify three key events corresponding to the three major plot points (#2 in Figure 3) from all of the edge events in a storyline and save them for later use in the prompt generation.
3.3 Generating Meta-Prompts
Although an ideal story should show and not tell, we observed that the LLM is prone to following instructions too explicitly. For example, the prompt of “the alternate storyline should contain new challenges…” would lead to events such as “Tony faces a new challenge of…”. Thus, we adopted a meta-prompting approach (i.e., prompting to generate prompts) which has shown to be an effective method to elicit better responses from LLMs de Wynter et al. (2023); Suzgun and Kalai (2024).
Given a storyline and every alternate decision () in each node, we can generate prompts that are tailored to every with concrete guiding questions using the key events we extracted from the previous phase.
Specifically, we ask GPT-4 to generate a prompt for each node, providing: (1) an alternate storyline that branches at event if the main character makes decision instead of ; (2) 5 concrete, guiding questions that expand: (2a) how would change and replace the Key Events & (2b) how would the main character make key decisions that overcome new challenges and propel the story forward; and (3) an ideal alternate storyline. Each prompt must also: (4) strictly adhere to the given events of the original storyline; and (5) be outputted as a list of events with as the first event. The tables in Appendix A show examples of the generated prompts.
By following the guiding questions with key events from the original storyline as context, GPT-4 is able to generate coherent stories that maintain a three-act story structure, guaranteeing steady narrative pacing and character development.
3.4 Branching & Merging the Tree
With meta-prompts generated for each node of the original storyline, the system recursively branches out at each node and writes the alternate storylines accordingly:
1.
Extract events from all edges of the original storyline graph and include them in the prompt for the corresponding node.
2.
Generate a list of new events that continue from the previous (extracted) events. (#4 in Figure 3)
3.
Create a branch from the list of new events (#5).
4.
Merge the branches of the previous events (the old plot tree) and new events to form a new branching plot tree (#6).
5.
Extract key events from the new storyline.
6.
Generate prompts for all the nodes on the branch using the procedure from Section 3.3.
7.
Recursively branch out on all the nodes of the branch and write alternate storylines. Return when all meta-prompts are used.
3.5 Narrating the Game
After our binary branching plot tree is complete, we have possible story endings in our game, with being the number of nodes/events in the original storyline. WHAT-IF traverses the tree to generate the narration for each node by prompting GPT-4 to narrate all events of the edge pointing towards the node from the previous time step. After that, it would narrate the state and goal of the character. Lastly, it provides two decisions and for the player to choose from. After the player makes a decision, WHAT-IF continues narration using the corresponding node in the tree. Lastly, we organize all narration into an Ink file—a scripting language developed by Inkle333https://www.inklestudios.com/ for writing interactive narratives—and use InkJS444https://github.com/y-lohse/inkjs to compile and generate the game online.
4 Qualitative Analysis & Discussion
To evaluate the effectiveness of our system, we use the plot of Iron Man taken from IMDB as the original story and compare the alternate storylines generated by GPT-4 using vanilla prompts and WHAT-IF prompts. Examples of the vanilla and WHAT-IF prompts and their corresponding output can be found in Appendix B.
4.1 Improving Branch Choices
As the baseline, vanilla prompting converts a story plot into a list of events and uses each event as potential branching points with a simple prompt of “Given the original storyline, imagine an alternate storyline where at event n, the main character makes decision x instead of y” to generate stories. The resulting stories suffer from the following problems:
(1) Non-choices. Since vanilla prompting considers every event as a potential branching point, any events that don’t involve decisions from the main character produce illogical branching events such as “Stark decides to get ambushed by the Ten Rings”. Although this aligns with the original event “Stark is critically wounded and captured by terrorists who demand he build a Jericho missile for them.”, this is not a decision Tony can make.
(2) Illogical decisions. Certain alternate decisions being proposed are impossible to consider given the circumstances and/or the character’s personality. For example, after getting captured by the Ten Rings, Tony Stark would never choose to stay in captivity.
(3) Retreat decisions. In interactive fiction writing, there is often a retreat option where the player decides to stay out of a situation. However, this would not be an ideal branching point that is meaningful to the story as it does not move the story forward. For example, “Tony decides not to save the villagers.
These decisions are attributed to branching based on events, thus the decisions are not always meaningful. WHAT-IF uses states and goals to capture major obstacles faced by the main character and considers the key decisions in response to the circumstances as potential branching points. This approach pushes the system to provide meaningful decisions for the player which will change the course of the story.
For example, when Tony Stark is imprisoned by Ten Rings, his state is “in captivity with a life-threatening injury”, and his goal is “to survive and escape captivity”. As an alternate decision is defined as a decision that the character could have made given the same state and goal, we avoid non-choices and illogical decisions by default. Since favoring a change in state forces the character to make a decision, the retreat decisions are eliminated.
4.2 Maintaining Story Structure
Story stagnation is the problem in fiction writing where the plot does not move forward and ends up losing the reader’s interest. We observed this phenomenon in some stories from the vanilla prompt’s tendency to hyper-focus on the alternate decision. For example, when Tony chooses to make a strategic plan to escape without building his famous Iron Man suit, the rest of the story turns into a journey of escaping from the cave and each event is a tiny step of the plan, with the story ending when Tony escapes and returns home.
In comparison, WHAT-IF uses the three-act structure to automatically extract the following three major plot points from Iron Man:
Inciting Incident: “Tony is critically wounded in an ambush by terrorists using Stark Industries weapons”
Crisis: “Stane steals Stark’s arc reactor, leaving him to die.”
Climax: “Stark survives using his original reactor, and battles Stane at Stark Industries”
These three key events are essential in creating the story arc of the original story, and the alternate stories should have similar events that build up tension, challenge the main character, and eventually come to a resolution. We include these events in the meta-prompt generation phase and ask GPT-4 to provide concrete guiding questions that expand on how the generated alternate decision is changed or replaced. With this prompt, for the alternate decision of “building the missile for Ten Rings hoping for escape later”, GPT-4 generated:
1. How would this alternate decision impact Stark’s relationship with Yinsen and their plan for escape?
2. What consequences would arise from the terrorists obtaining a functioning Jericho missile?
3. How would Stark manage to escape without the assistance of a suit of armor?
4. In what ways would Stark confront and expose Stane’s dealings without the identity of Iron Man?
5. How would Stark’s decision affect his ideology and future decisions regarding weapon manufacturing and his role in global conflicts?
By answering these concrete guiding questions specifically related to major plot points from the original story, GPT-4 ensures that (1) the consequence of the alternate questions are extensively explored, (2) the unaffected original events still occur, e.g. Stane’s betrayal, and (3) the changes to the story arc are considered.
After new events are generated, they are organized into a branching plot tree again. A new set of key events is extracted from the merged tree and fed to generate more context-specific prompts. We make sure to only include the key events that have not yet occurred at the branching event to prevent changing the past.
4.3 Controllable LLM outputs
While text-adventure games where any action could potentially cause the player’s demise, many modern interactive fictions prefer to balance branching narratives, where each branch has approximately the same length to prevent the story from ending too quickly. WHAT-IF builds a full binary branching plot tree to ensure complete balance. Furthermore, GPT-4 is told to keep the generated alternate plots to the same number of events as the original plot. To control GPT-4’s output, we use the JSON mode with predefined JSON schema for all prompting functions to ensure the correct number of events or nodes are generated. Examples of the JSON schema are included in Appendix C.
5 Conclusion
In this paper, we presented an interactive fiction system called WHAT-IF, which uses LLMs to generate alternate timelines from pre-written storylines. WHAT-IF focuses on character-driven generation to create engaging branching points in the narratives, uses meta-prompting to consider the narrative from a broader perspective using the three-act story structure, and maintains consistent storytelling using the meta-prompts and an external graph structure. We hope to see future story generation systems use meta-prompting to encourage controlled, long-form generation.
Limitations
Two of the main limitations of using our system are time and money. Branching takes about a minute to generate each branch. We opted to generate the whole tree ahead of time—which saves time while playing—at the expense of waiting for longer to start.
GPT-4 is priced by how many tokens it has used as input and how many tokens it has generated. This makes any widespread use of a system that relies on GPT-4 to be impractical. If we were to release WHAT-IF to the general public, it would need to be converted to a free-to-use LLM of similar ability. WHAT-IF might need to be converted to a few-shot system, but the overall structure and prompting techniques we used will most likely still help the LLM.
Also, as with all large language models, the generated text cannot be predicted ahead of time, which can lead to potentially harmful stories.
Lastly, our experiments with WHAT-IF were in English only. Because of its architecture, WHAT-IF is bound to the LLM underneath it. As most LLMs work best with English at the moment, this limits our system as well.
Appendix B Prompt Examples
B.1 Vanilla Prompt
Extract Events from Plot:
Get decisions from each event:
Write new storylines:
B.2 WHAT-IF Prompts
The following are all of the prompts we use for the system, in order but without the JSON schema we ask it to use. The JSON schemas for each prompt can be found in Appendix C and would be filled in via the JSON_SCHEMA variable. These are the prompts for generating the tree (plot-to-tree), extracting key events, generating meta-prompts, writing a new storyline, and generating the final game narration, respectively.
Appendix C JSON Schema
The following are all of the JSON that are used in the prompts above. They are for generating the tree (plot-to-tree), extracting key events, generating meta-prompts, writing a new storyline, generating the final game narration.